# DualContrast: Unsupervised Disentangling of Content and Transformations with Implicit Parameterization

### v1

@@ 3.3 Variational Inference of Content and Transformation Factors
The log-likelihood term is maximized using a reconstruction loss $L_{\text{recon}}$. The prior for $\mathbf{c}$ and $\mathbf{z}$, $p(\mathbf{c})$ and $p(\mathbf{z})$ respectively are assumed to be standard normal distribution. In short, we write the KL terms as $L_{\text{KL}(c)}$ and $L_{\text{KL}(z)}$. To regularize the effect of prior on the final objective separately for two latent factors, we scale the two losses with two hyperparameters $\gamma_{c}$ and $\gamma_{z}$. Our loss function for variational inference becomes:
@@
We observe that by deactivating the KL term associated with any one of the factors, the likelihood tends to be optimized through the other latent factor. For instance, if $\gamma_{c}$ is set as 00 and $\gamma_{z}$ has a large value, then the $L_{\text{rec}}$ is optimized through $\mathbf{c}$ only. As a consequence, $\mathbf{c}$ captures all the information in $\mathbf{x}$, whereas $\mathbf{z}$ becomes totally independent of the data. In such a scenario, $\mathbf{z}$ will be a random factor with no relation to the transformation factor of the data, which is not desirable.
3.4 Creating Contrastive Pair with respect to Content Latent Factor
@@
We adopt a contrastive strategy for content-transformation disentanglement. We first develop a strategy to impose a structure to the content latent factor with contrastive learning. To this end, we create negative and positive pairs for each data sample w.r.t. the content factor. Hence, we randomly pick two samples $\mathbf{x}$ and $\mathbf{x}_{\text{neg}(\mathbf{c})}$, where $\mathbf{x},\mathbf{x}_{\text{neg}(\mathbf{c})}\in\mathcal{X}$ and regard them as negative pairs of each other w.r.t the content factor. Such strategy is commonly used in traditional self-supervised learning approaches [16, 33]. Considering the high heterogeneity present in the dataset, evidenced by a large number of classes, this approach accounts for the likelihood that any two randomly selected samples will have different content. To generate positive pairs w.r.t. the content factor, we pick any sample $\mathbf{x}$ and transform it with a transformation $T\in\mathcal{T}$ to generate $\mathbf{x}_{\text{pos}(\mathbf{c})}$. We use the encoder $q_{\psi}$ to generate content factors $\mathbf{c}$, $\mathbf{c}_{\text{pos}}$, and $\mathbf{c}_{\text{neg}}$ respectively from $\mathbf{x}$, $\mathbf{x}_{\text{pos}(\mathbf{c})}$, and $\mathbf{x}_{\text{neg}(\mathbf{c})}$. We use a contrastive loss to penalize the distance between $\mathbf{c}$ and $\mathbf{c}_{\text{pos}}$ and the similarity between $\mathbf{c}$ and $\mathbf{c}_{\text{neg}}$. Our contrastive loss for content latent factor can be written as:
$L_{\text{con}(\mathbf{c})}=L_{\text{dist}}(\mathbf{c},\mathbf{c}_{\text{pos}})+L_{\text{sim}}(\mathbf{c},\mathbf{c}_{\text{neg}})$ (3)
To implement $L_{\text{dist}}$ and $L_{\text{sim}}$, we use mean absolute cosine distance and mean absolute cosine similarity, respectively.
3.5 Creating Contrastive Pair with respect to Transformation Latent Factor
@@
Imposing contrastive loss on the content latent factor satisfies condition 1 of the content-transformation disentanglement (Section 3.1). However, the content-transformation disentanglement remains only partial since it does not satisfy condition 2. To solve this issue, we design another contrastive loss w.r.t. the transformation latent factor. Designing a contrastive loss that explicitly enforces condition 2 is not possible, given the transformation that may equalize the transformation factor of two samples can not be approximated. To this end, we design a contrastive loss that implicitly encourages condition 2. We validated the design experimentally in Section 4.
@@
The negative pair of a sample $\mathbf{x}\in\mathcal{X}$ w.r.t. transformation latent factor is generated while creating a positive pair of it w.r.t content factor. This is because a transformation $T$ on $\mathbf{x}$ alters its transformation, but not its contents. So, we regard $\mathbf{x}_{\text{pos}(\mathbf{c})}$ as $\mathbf{x}_{\text{neg}(\mathbf{z})}$. However, creating positive pairs of samples w.r.t. transformation is very challenging as they can not be created in the input space, unlike others. To this end, we adopted a strategy of creating pairs of samples from the same distribution in the transformation factor embedding space. We randomly sample $\mathbf{z}^{(1)}$ and $\mathbf{z}^{(2)}$ from $\mathcal{N}(0,1)$ and use the decoder $p_{\theta}$ to generate samples $\mathbf{x}^{(1)}_{\text{pos}(\mathbf{z})}$ and $\mathbf{x}^{(2)}_{\text{pos}(\mathbf{z})}$ from them (Fig. 1 (c)). In the decoder, as content input, we use content factor $\mathbf{c}^{(1)}$ and $\mathbf{c}^{(2)}$ obtained by encoding two random samples $\mathbf{x}^{(1)}$ and $\mathbf{x}^{(2)}$. Consequently, $\mathbf{x}^{(1)}_{\text{pos}(\mathbf{z})}=p_{\theta}(\mathbf{c}^{(1)},\mathbf{z}^{(1)})$ and $\mathbf{x}^{(2)}_{\text{pos}(\mathbf{z})}=p_{\theta}(\mathbf{c}^{(2)},\mathbf{z}^{(2)})$ serve as the positive pairs w.r.t. transformation latent factor. We then use the encoder $q_{\psi}$ to generate $\mathbf{z}$, $\mathbf{z}_{\text{neg}}$, $\mathbf{z}^{(1)}_{\text{pos}}$, and $\mathbf{z}^{(2)}_{\text{pos}}$ from $\mathbf{x}$, $\mathbf{x}_{\text{neg}(\mathbf{z})}$, $\mathbf{x}^{(1)}_{\text{pos}(\mathbf{z})}$, and $\mathbf{x}^{(2)}_{\text{pos}(\mathbf{z})}$ respectively. We use a contrastive loss to penalize the distance between $\mathbf{z}^{(1)}_{\text{pos}}$ and $\mathbf{z}^{(2)}_{\text{pos}}$ and similarity between $\mathbf{z}$ and $\mathbf{z}_{\text{neg}}$. Our contrastive loss for transformation latent factor can be written as:
$L_{\text{con(z)}}=L_{\text{dist}}(\mathbf{z}^{(1)}_{\text{pos}},\mathbf{z}^{(2)}_{\text{pos}})+L_{\text{sim}}(\mathbf{z},\mathbf{z}_{\text{neg}})$ (4)
Similar to $L_{\text{con}(\mathbf{c})}$, we use mean absolute cosine distance and mean absolute cosine similarity to implement $L_{\text{dist}}$ and $L_{\text{sim}}$, respectively.
@@ 4 Experiments & Results
Table 1: Quantitative Results of unsupervised Content-Transformation Disentangling Methods on MNIST and LineMod datasets. This evaluation only shows the disentanglement performance w.r.t ground truth content factors. We report mean $\pm$ std. dev. over model training by setting 3 different random seeds. Here, for Harmony and SpatialVAE, the transformation latent factor is restricted to dimension 3. For the other methods, there is no such restriction on the transformation latent factor. For SimCLR, (Disc) means discriminative model (without any reconstruction), and (Gen) means Generative model (with reconstruction).
MNIST LineMod
Method $D_{\text{score}}(c|c)(\uparrow)$ $D_{\text{score}}(c|s)(\downarrow)$ $D_{\text{score}}(c|c)(\uparrow)$ $D_{\text{score}}(c|s)(\downarrow)$
SpatialVAE [2] 0.81 $\pm$ 0.01 0.28 $\pm$ 0.02 0.95 $\pm$ 0.02 0.32 $\pm$ 0.01
Harmony [4] 0.82 $\pm$ 0.01 0.31 $\pm$ 0.02 0.90 $\pm$ 0.01 0.56 $\pm$ 0.02
SimCLR (Disc) with rotation [15] 0.58 $\pm$ 0.01 0.60 $\pm$ 0.01 0.62 $\pm$ 0.01 0.40 $\pm$ 0.01
SimCLR (Gen) with rotation [15] 0.53 $\pm$ 0.01 0.67 $\pm$ 0.02 0.61 $\pm$ 0.02 0.79 $\pm$ 0.01
VAE with two latent space 0.63 $\pm$ 0.00 0.63 $\pm$ 0.00 0.87 $\pm$ 0.00 0.87 $\pm$ 0.00
VITAE [1] 0.77 $\pm$ 0.00 0.32 $\pm$ 0.00 0.92 $\pm$ 0.00 0.90 $\pm$ 0.00
DualContrast (w/o $L_{\text{con(c)}}$) 0.87 $\pm$ 0.02 0.21 $\pm$ 0.02 0.86 $\pm$ 0.01 0.31 $\pm$ 0.02
DualContrast (w/o $L_{\text{con(z)}}$) 0.79 $\pm$ 0.04 0.85 $\pm$ 0.02 0.79 $\pm$ 0.03 0.86 $\pm$ 0.03
DualContrast 0.89 $\pm$ 0.01 0.31 $\pm$ 0.02 0.95 $\pm$ 0.01 0.48 $\pm$ 0.02
@@
Benchmark Datasets: We report results on MNIST [34], LineMod [35], Starmen [36], and a realistically simulated protein subtomogram dataset. Only LineMod features RGB images among these datasets, while the other three consist of grayscale images. The subtomogram is a 3D volumetric dataset, whereas the others are 2D images.
@@
Baselines: We used the unsupervised content-transformation disentanglement methods, SpatialVAE [2], Harmony [4], VITAE [1] with explicit parameterization and InfoNCE loss based standard self-supervised learning method with rotation augmentation [15] as baseline approaches. Several generic disentangled representation learning methods, such as beta-VAE, Factor-VAE, beta-TCVAE, DIP-VAE, etc., are available that do not aim to disentangle specific factors such as content or transformation in the data but rather aim to disentangle all factors of variation. However, previous studies [2, 1] have shown that such methods perform worse than methods specifically aiming to disentangle content and transformation. Therefore, we did not use these methods as baselines in our experiments.
@@
Implementation Details: We implemented the encoder using a Convolutional Neural Network (CNN) and the decoder using a Fully Connected Network (FCN). Our experiments used the same latent dimension for content and transformation factors, except for Harmony [4] and SpatialVAE [2], where a specific dimension needs to be used for transformation latent factors. We train our models for $200$ epochs with a learning rate of $0.0001$ in NVIDIA RTX A500 GPUs and AMD Raedon GPUs. We optimize the model parameters with Adam optimizer.
Evalutation Metrics: We assess DualContrast based on two criteria: (1) the quantitative disentanglement of content and transformation factors and (2) the qualitative capability of image generation through content-transformation transfer.
Figure 2: Qualitative content-transformation transfer based image generation results for MNIST. (a), (b), (c), and (d) shows the results with Harmony [4], Spatial-VAE [2], VITAE [1], and DualContrast respectively. Erroneous generations (both in terms of content and transformation) are marked within red boxes.
@@
For (1), there exists several metrics, e.g., MIG score, $D_{\text{score}}$, etc. Locatello et al [37] demonstrated that these metrics are highly correlated. To this end, similar to [1], we only use $D_{\text{score}}$ to measure the disentanglement. $D_{\text{score}}$ is simply a measurement of how well the ground truth factors can be predicted from the corresponding latent factors. In our scenario, there are four such quantities - (a) predictivity of content labels given content factor $D_{\text{score}}(c|c)$ or transformation factor $D_{\text{score}}(c|s)$ (b) predictivity of transformation labels given content factor $D_{\text{score}}(s|c)$ or transformation factor $D_{\text{score}}(s|s)$. For MNIST and LineMod, we do not have any transformation labels, so we only reported values for (a). Similar to [15], we use linear Logistic Regression Classifiers to measure predictivity ($D_{\text{score}}$) across all our experiments.
Figure 3: tSNE embedding plots of content latent factor learned by the unsupervised content-style disentanglement methods. (a), (b), (c), and (d) shows the results for Harmony [4], Spatial-VAE [2], C-VITAE [1], and DualContrast respectively. Overall, DualContrast shows superior performance.
The quantitative measurement, particularly when transformation labels are not available, might be misleading. To this end, we perform qualitative evaluation by generating images while transferring content and transformation from original images from the dataset. We take a set of $K$ images $\mathbf{x}$ and create a $K\times K$ grid of $K^{2}$ images using the different combinations of content and transformation factors from those $K$ images. In the grid, we keep the content factor and transformation factor consistent across the columns and the rows respectively. In other words, the image at position $(i,j)$ in the grid is generated using the content of image $x_{j}$ and transformation of image $x_{i}$. For the protein subtomogram dataset, we also provide qualitative results through embedding visualization (Figure 8).
@@
4.1 DualContrast disentangles writing style from MNIST handwritten digit images
Similar to [4, 2, 1], we start our experiments with the MNIST dataset of handwritten digits. We performed a quantitative comparison between DualContrast and the baseline approaches, as shown in Table 1. We observed that DualContrast achieves higher $D_{score}(c|c)$ compared to other explicitly parameterized approaches [4, 2, 1].

### v2

@@ 3.3 Variational Inference of Content and Transformation Codes
The log-likelihood term is maximized using a reconstruction loss $L_{\text{recon}}$. The prior for $\mathbf{c}$ and $\mathbf{z}$, $p(\mathbf{c})$ and $p(\mathbf{z})$ respectively are assumed to be standard normal distribution. In short, we write the KL terms as $L_{\text{KL}(c)}$ and $L_{\text{KL}(z)}$. To regularize the effect of prior on the final objective separately for two latent codes, we scale the two losses with two hyperparameters $\gamma_{c}$ and $\gamma_{z}$. Our loss function for variational inference becomes:
@@
We observe that by deactivating the KL term associated with any one of the codes, the likelihood tends to be optimized through the other latent code. For instance, if $\gamma_{c}$ is set as 00 and $\gamma_{z}$ has a large value, then the $L_{\text{rec}}$ is optimized through $\mathbf{c}$ only. As a consequence, $\mathbf{c}$ captures all the information in $\mathbf{x}$, whereas $\mathbf{z}$ becomes totally independent of the data. In such a scenario, $\mathbf{z}$ will be a random code with no relation to the transformation of the data, which is not desirable.
3.4 Creating Contrastive Pair with respect to Content
@@
We adopt a contrastive strategy for c-z disentanglement. We first develop a strategy to impose a structure to the content latent code with contrastive learning. To this end, we create negative and positive pairs for each data sample w.r.t. the content code. Hence, we randomly pick two samples $\mathbf{x}$ and $\mathbf{x}_{\text{neg}(\mathbf{c})}$, where $\mathbf{x},\mathbf{x}_{\text{neg}(\mathbf{c})}\in\mathcal{X}$ and regard them as negative pairs of each other w.r.t the content code. Such strategy is commonly used in traditional self-supervised learning approaches [16, 24]. Considering the high heterogeneity present in the dataset, evidenced by a large number of classes, this approach accounts for the likelihood that any two randomly selected samples will have different content. To generate positive pairs w.r.t. the content code, we pick any sample $\mathbf{x}$ and transform it with a transformation $T\in\mathcal{T}$ to generate $\mathbf{x}_{\text{pos}(\mathbf{c})}$. We use the encoder $q_{\psi}$ to generate content codes $\mathbf{c}$, $\mathbf{c}_{\text{pos}}$, and $\mathbf{c}_{\text{neg}}$ respectively from $\mathbf{x}$, $\mathbf{x}_{\text{pos}(\mathbf{c})}$, and $\mathbf{x}_{\text{neg}(\mathbf{c})}$. We use a contrastive loss to penalize the distance between $\mathbf{c}$ and $\mathbf{c}_{\text{pos}}$ and the similarity between $\mathbf{c}$ and $\mathbf{c}_{\text{neg}}$. Our contrastive loss for content latent code can be written as: $L_{\text{con}(\mathbf{c})}=L_{\text{dist}}(\mathbf{c},\mathbf{c}_{\text{pos}})+L_{\text{sim}}(\mathbf{c},\mathbf{c}_{\text{neg}})$. To implement $L_{\text{dist}}$ and $L_{\text{sim}}$, we use mean absolute cosine distance and mean absolute cosine similarity, respectively.
3.5 Creating Contrastive Pair with respect to Transformation
@@
Imposing contrastive loss on the content code satisfies condition 1 of the c-z disentanglement (Section 3.1). However, the c-z disentanglement remains only partial since it does not satisfy condition 2. To solve this issue, we design another contrastive loss w.r.t. the transformation code. Designing a contrastive loss that explicitly enforces condition 2 is not possible, given the transformation code that may equalize the transformation of two samples can not be approximated. To this end, we design a contrastive loss that implicitly encourages condition 2. We validated the design experimentally in Section 4.
@@
The negative pair of a sample $\mathbf{x}\in\mathcal{X}$ w.r.t. transformation is generated while creating a positive pair of it w.r.t content. This is because a transformation $T$ on $\mathbf{x}$ alters its transformation, but not its contents. So, we regard $\mathbf{x}_{\text{pos}(\mathbf{c})}$ as $\mathbf{x}_{\text{neg}(\mathbf{z})}$. However, creating positive pairs of samples w.r.t. transformation is very challenging as they can not be created in the input space, unlike others. To this end, we adopted a strategy of creating pairs of samples from the same distribution in the transformation embedding space. We randomly sample $\mathbf{z}^{(1)}$ and $\mathbf{z}^{(2)}$ from $\mathcal{N}(0,1)$ and use the decoder $p_{\theta}$ to generate samples $\mathbf{x}^{(1)}_{\text{pos}(\mathbf{z})}$ and $\mathbf{x}^{(2)}_{\text{pos}(\mathbf{z})}$ from them (Fig. 2 (c)). In the decoder, as content input, we use content codes $\mathbf{c}^{(1)}$ and $\mathbf{c}^{(2)}$ obtained by encoding two random samples $\mathbf{x}^{(1)}$ and $\mathbf{x}^{(2)}$. Consequently, $\mathbf{x}^{(1)}_{\text{pos}(\mathbf{z})}=p_{\theta}(\mathbf{c}^{(1)},\mathbf{z}^{(1)})$ and $\mathbf{x}^{(2)}_{\text{pos}(\mathbf{z})}=p_{\theta}(\mathbf{c}^{(2)},\mathbf{z}^{(2)})$ serve as the positive pairs w.r.t. transformation. We then use the encoder $q_{\psi}$ to generate $\mathbf{z}$, $\mathbf{z}_{\text{neg}}$, $\mathbf{z}^{(1)}_{\text{pos}}$, and $\mathbf{z}^{(2)}_{\text{pos}}$ from $\mathbf{x}$, $\mathbf{x}_{\text{neg}(\mathbf{z})}$, $\mathbf{x}^{(1)}_{\text{pos}(\mathbf{z})}$, and $\mathbf{x}^{(2)}_{\text{pos}(\mathbf{z})}$ respectively. We use a contrastive loss to penalize the distance between $\mathbf{z}^{(1)}_{\text{pos}}$ and $\mathbf{z}^{(2)}_{\text{pos}}$ and similarity between $\mathbf{z}$ and $\mathbf{z}_{\text{neg}}$. Our contrastive loss for transformation can be written as: $L_{\text{con(z)}}=L_{\text{dist}}(\mathbf{z}^{(1)}_{\text{pos}},\mathbf{z}^{(2)}_{\text{pos}})+L_{\text{sim}}(\mathbf{z},\mathbf{z}_{\text{neg}})$. Similar to $L_{\text{con}(\mathbf{c})}$, we use mean absolute cosine distance and mean absolute cosine similarity to implement $L_{\text{dist}}$ and $L_{\text{sim}}$, respectively.
Why rotation?: We used rotation as a transformation to create contrastive pairs because only rotation provided generalization to other shape-based transformations in the dataset, compared to other transformations such as translation, scale, blur, or color-based transformations (Details in Appendix A.2.3).
@@
We optimize $L_{\text{vae}}$ for both $\mathbf{x}$ and the transformed image $\mathbf{x}_{\text{pos}(\mathbf{c})}$ (same as $\mathbf{x}_{\text{neg}(\mathbf{z})}$), which yielded better experimental result.
@@ 4 Experiments & Results
Table 1: Quantitative Results of unsupervised Content-Transformation Disentangling Methods. Among these methods, SpatialVAE and Harmony put constraints on the $\mathbf{z}$ code dimension given their explicit parameterization, others do not. The std. dev. over model training by setting 3 different random seeds remains within ±$\pm 0.04$. $D_{\text{score}}$ is written as $D$. For $D(c|c)$ and $D(z|z)$, higher is better. For $D(c|z)$ and $D(z|c)$, lower is better. For $SAP(c)$ and $SAP(z)$, higher is better.
MNIST LineMod Protein Subtomogram
Method $D(c|c)$ $D(c|z)$ $SAP(c)$ $D(c|c)$ $D(c|z)$ $SAP(c)$ $D(c|c)$ $D(c|z)$ $SAP(c)$ $D(z|z)$ $D(z|c)$ $SAP(z)$
SpatialVAE 0.81 0.28 0.53 0.95 0.32 0.63 0.69 0.93 0.24 0.71 0.57 0.14
Harmony 0.82 0.31 0.51 0.90 0.56 0.34 0.95 0.01 0.94 0.52 0.90 0.38
SimCLR (Discriminative) 0.58 0.60 0.02 0.62 0.40 0.22 0.39 0.69 0.30 0.51 0.52 0.01
SimCLR (Generative) 0.53 0.67 0.14 0.61 0.79 0.18 0.53 0.59 0.06 0.55 0.54 0.01
VAE with $2$ latent space 0.63 0.63 0.0 0.87 0.87 0.0 0.79 0.82 0.03 0.85 0.85 0.0
VITAE 0.77 0.32 0.45 0.92 0.90 0.02 - - - - - -
DualContrast (w/o $L_{\text{con(c)}}$) 0.87 0.21 0.66 0.86 0.31 0.55 0.93 0.81 0.13 0.78 0.81 0.03
DualContrast (w/o $L_{\text{con(z)}}$) 0.79 0.85 0.06 0.79 0.86 0.07 0.99 0.86 0.13 0.86 0.63 0.23
DualContrast 0.89 0.31 0.58 0.95 0.48 0.47 1.00 0.56 0.44 0.86 0.64 0.22
@@
Benchmark Datasets: We report results on MNIST [25], LineMod [26], and a realistically simulated protein subtomogram dataset. Only LineMod features RGB images among these datasets, while the others consist of grayscale images. The subtomogram is a 3D volumetric dataset, whereas the others are 2D images. For a detailed discussion of the datasets, we refer to the Appendix.
@@
Baselines: We used the unsupervised c-z disentanglement methods, SpatialVAE [2], Harmony [4], VITAE [1] with explicit parameterization and InfoNCE loss based standard self-supervised learning method with rotation augmentation [15] as baseline approaches. Several generic disentangled representation learning methods, such as beta-VAE, Factor-VAE, beta-TCVAE, DIP-VAE, etc., are available that do not aim to disentangle specific factors such as content or transformation in the data but rather aim to disentangle all factors of variation. However, previous studies [2, 1] have shown that such methods perform worse than methods specifically aiming to disentangle content and transformation. Therefore, we did not use these methods as baselines in our experiments.
@@
Implementation Details: We implemented the encoder using a Convolutional Neural Network (CNN) and the decoder using a Fully Connected Network (FCN). Our experiments used the same latent dimension for content and transformation codes, except for Harmony [4] and SpatialVAE [2], where a specific dimension needs to be used for transformation codes. We train our models for $200$ epochs with a learning rate of $0.0001$ in NVIDIA RTX A500 GPUs and AMD Raedon GPUs. We optimize the model parameters with Adam optimizer.
The hyperparameters $\gamma_{z}$ and $\gamma_{c}$ controls how close the z and c distribution will be to $\mathcal{N}(0,I)$. We found setting a small value $\approx 0.01$ for both provides optimal results in our experiments. Further details of the overall implementation are provided in the Appendix A.3.
Evalutation: We evaluate the methods based on two criteria: (1) the informativeness of the predicted $\mathbf{c}$ and $\mathbf{z}$ codes w.r.t the ground truth $\mathbf{c}$ and $\mathbf{z}$ factors, (2) the separateness of of the predicted $\mathbf{c}$ and $\mathbf{z}$ codes. We performed these evaluations both quantitatively and qualitatively.
@@
For quantitative evaluations, there exists several metrics, e.g. , MIG score, $D_{\text{score}}$, SAP score, etc. [27] demonstrated that these metrics are highly correlated. To this end, we only used $D_{\text{score}}$ and SAP score to measure the disentanglement. $D_{\text{score}}$ is simply a measurement of how well the ground truth (gt) factors can be predicted from the corresponding latent codes. In our scenario, there are four such quantities - (a) predictivity of content gt given $\mathbf{c}$ code $D_{\text{score}}(c|c)$ or $\mathbf{z}$ code $D_{\text{score}}(c|z)$ (b) predictivity of transformation gt given $\mathbf{c}$ code $D_{\text{score}}(z|c)$ or $\mathbf{z}$ code $D_{\text{score}}(z|z)$. For MNIST and LineMod, we do not have any transformation gt, so we only reported values for (a). Similar to [15], we use linear Logistic Regression Classifiers to measure predictivity ($D_{\text{score}}$). The higher the $D_{\text{score}}$, the more informative is the latent code w.r.t. the gt. SAP score for a gt factor is defined as the difference between the highest and second highest predictivities for it given any latent codes. In our case, SAP(c) can be defined as $|D_{\text{score}}(c|c)-D_{\text{score}}(c|z)|$ and SAP(z) as $|D_{\text{score}}(z|z)-D_{\text{score}}(z|c)|$. The higher the SAP score, higher the separateness of the latent codes. For qualitative evaluations, we performed downstream tasks like generating images through unsupervised content-transformer transfer and latent space visualization. For protein subtomograms, we performed subtomogram averaging [28] to identify proteins resulting from the latent space clusters.
@@
4.1 DualContrast disentangles several writing styles from MNIST images